OpenStudy (anonymous):

In an experiment to study the growth of bacteria, a medical student measured 5000 bacteria at t=0 minutes and 8000 at t=10 minutes. Assuming the growth of bacteria is exponential, how many bacteria are present at t=30 minutes.

OpenStudy (anonymous):

I'm having a hard time finding constant 'k'. Or modeling it in general...

OpenStudy (amistre64):

A= Pe^rt A/P = e^rt ln(A/P) = rt ln(A/P)/t = r

OpenStudy (anonymous):

one sec.

OpenStudy (amistre64):

ln(8000/5000)/10 = r

OpenStudy (anonymous):

Oh, I DO know P. I didn't think I did, but it's just 5000. Thanks!

OpenStudy (amistre64):

yep